twice one number added to three times another gives 21. find the numbers, if the difference between them is 3.
Answer by elima(1433)

twice one number added to three times another gives 21. find the numbers, if the difference between them is 3.
2x+3y=21
x-y=3
Lets use substitution;
first change the second equation so you have what x equals;
x=3+y
Now plug that into the first equation for x;
2(3+y)+3y=21
6+2y+3y=21
5y=21-6
5y=15
y=3
Now plug your y value into the second equaiton;
x-3=3
x=3+3
x=6
check;
6-3=3
2(6)+3(3)=21
12+9=21
:)